OpenStudy (anonymous):

simplify : 1/3 + ( - 2/3) + 5/6

Parth (parthkohli):

= -2/6 + 5/6 = 3/6 = 1/2

OpenStudy (anonymous):

1/3 + (-2/3) = - 1/3 + 5/6 = 4/6 which is 2/3 simplified.

Parth (parthkohli):

1/2 is the answer ^^ Dr. dave

OpenStudy (anonymous):

Indeed, since 1/3 = 2/6